OpenStudy (anonymous):

find the limit as x approachs 0: 8 sin^2 (x) / 1- cos(x)

OpenStudy (agent47):

rewrite sin^2(x) as 1-Cos^2(x)

OpenStudy (anonymous):

Hint: sin^2 x= 1-cos^2 x = (1+cosx) (1-cosx)

OpenStudy (agent47):

That can then be factored to (1-Cosx)(1+Cosx)

OpenStudy (agent47):

Then the bottom cancels out, on top you have 1+Cosx 8(1+Cosx)=8(1+1)=16

OpenStudy (anonymous):

cool. could you use l'hospitals rule too?